OpenStudy (anonymous):

A turtle and a hare race. The hare takes off at 4.5 m/s for 15 minutes and the turtle takes off at 0.45 m/s. The hare then slows down to 1.5 m/s for 10 mins. Then it falls asleep for 20 minutes. It wakes up and sprints for the next 15 mins at 6.2 m/s. a. What is the hare's avg. velocity? b. What is the total displacement of the hare? c. What is the turtle's displacement if he continued moving at the same rate for the same amount of time as the hare's total trip?

OpenStudy (ramirocruzo):

Can you tell me whats the duration for which Hare ran with 4.5m/s????

OpenStudy (anonymous):

15 minutes, sorry.

OpenStudy (ramirocruzo):

\[v _{avg}=\frac{ Total Displacement }{Total Time }\] So, \[v _{avg}=2.925m/s\]for Hare Total displacement=10530m {Vel.xTime(In sec)} \[Turtle's Displacement=Vel. \times Time=0.45 \times 3600= 16202m\]